If rounding required round to the 4th decimal point.

2. 0.8/2 points I Previous Answers PSE6 5.P.014 My Notes Three forces acting on an object are given by F1 (1.90 i 2.40 j) N, F2 -(-4.60 İ ) N, and F3-( 46.51-42.0 j ) N. The object experiences an acceleration of magnitude 3.75 m/s2. (a) What is the direction of the acceleration? o (from the positive x axis) (b) What is the mass of the object? 15.745 (c) If the object is initially at rest, what is its speed after 11.0 s? 4125m/s (d) What are the velocity components of the object after 11.0 s? kg 43.80 1x Your response differs from the correct answer by more than 10%. Double check your calculations. i 39.6 Your response differs from the correct answer by more than 10%. Double check your calculations. j) m/s
Answer #1

Given is:-

Fı = (1.90i + 2.40)Ν.Ћ, = (-4.600 V. Fl = (46.52-42.0

Magnitude of acceleration is a=3.75 m/s2

Now,

part-a

The Net force is

Net = F1 + F2 + F3

by plugging all the values we get

Fyet (1.90-4.60 + 46.5)ị t (2.40 + 0-42.0)j

which gives us

FNet(43.8)i(-39.6)

The direction of the acceleration or force is given by

heta = tan^{-1}(rac{F_y}{F_x})

by plugging all the values we get

39.6 43.8 θ tan-1 (

θ 12.12

or

θ-360-42.12

θ 317.9° counter clockwise from the +x axis

Part-b

Magnitude of the Force

FI-(43.8)2(-39.6)2

which gives us

F59.05N

thus the mass of the object is

F 59.05 3.75

m 15.75kg

Part-c

Initial speed u=0 m/s

The final speed at t=11 seconds will be given by the first equation of motion thus

v = u + at

0 3.75(11)

which gives us

41.25m/s

Part-d

The final velocity vector is given by

?7m

by plugging all the values we get

7 -0[(43.87)-(39.61 15.75

(30.6m/s)i (27.66m/s)j